# A Survey on Hallucination in Large Vision-Language Models

### v1

@@ Abstract
Recent development of Large Vision-Language Models (LVLMs) has attracted growing attention within the AI landscape for its practical implementation potential. However, “hallucination”, or more specifically, the misalignment between factual visual content and corresponding textual generation, poses a significant challenge of utilizing LVLMs. In this comprehensive survey, we dissect LVLM-related hallucinations in an attempt to establish an overview and facilitate future mitigation. Our scrutiny starts with a clarification of the concept of hallucinations in LVLMs, presenting a variety of hallucination symptoms and highlighting the unique challenges inherent in LVLM hallucinations. Subsequently, we outline the benchmarks and methodologies tailored specifically for evaluating hallucinations unique to LVLMs. Additionally, we delve into an investigation of the root causes of these hallucinations, encompassing insights from the training data and model components. We also critically review existing methods for mitigating hallucinations. The open questions and future directions pertaining to hallucinations within LVLMs are discussed to conclude this survey.
@@ 1 Introduction
Existing studies Rohrbach et al. (2018); Li et al. (2023b); Hu et al. (2023); Zhai et al. (2023) have addressed the issue of hallucination in image captioning models with a predominant focus on the “presence of objects”, specifically whether the objects depicted in a given image are accurately described by the text generated by the model. In contrast to image captioning models trained within closed domains, LVLMs leverage the robust understanding and expressive capabilities of LLMs to obtain more detailed and explicable generated descriptions. However, these enhanced abilities also diversify and potentially exacerbate hallucination, which is not only limited to object existence, but also manifested in the descriptive errors such as attribute and relation errors. We focus on visual hallucinations, referring to all inconsistencies between the semantic content conveyed by the image and the textual content generated by the model.
@@
While the LLM community has extensively discussed the causes of hallucinations in LLMs, LVLMs’ visual modality introduces unique challenges in analyzing these occurrences. We conduct a thorough analysis of hallucinations in LVLMs, focusing on the training data and model characteristics. Our analysis indicates that hallucinations in LVLMs are caused not only by the generative nature of LLMs but also by biased training data, the inability of vision encoders to accurately ground images, misalignment among different modalities, insufficient context attention, and many other factors. Following this, we provide a comprehensive overview of the existing hallucination mitigation methods. Corresponding to the causes, current mitigation approaches predominantly focus on the optimization of training data, refinement of various modules within LVLMs, and post-processing of generated outputs. These methods are employed to diminish the occurrence of hallucinations, thereby yielding more faithful responses. In the end, we list several considerable directions for developing the research of hallucinations in LVLMs.
@@ 2.1 Large Vision-Language Models
LVLM architectures typically comprise three components: a visual encoder, a modality connection module, and a LLM. The visual encoder, often an adaption of the CLIP vision encoder Radford et al. (2021), transforms input images into visual tokens. The connection module is designed to align visual tokens with the word embedding space of the LLM, ensuring that the LLM can handle the visual information. There are various methods for modality alignment, including cross-attention Alayrac et al. (2022), adapters Gao et al. (2023), Q-Formers Li et al. (2023a); Dai et al. (2023a); Zhu et al. (2023), and simpler structures such as linear layers or multi-layer perceptrons (MLP) Liu et al. (2023c); Chen et al. (2023b); Liu et al. (2023b). The LLM functions like the central processing unit within LVLMs, receiving aligned visual and textual information and subsequently synthesizing this information to produce responses.
@@ 3 Evaluation Methods and Benchmarks
FAITHSCORE Min et al. (2023) , leaf, text width=15em ] ] [ Model-based
@@ 3.1 Evaluation on Non-Hallucinatory Generation
End-to-end methods directly assess the performance of LVLMs by evaluating their responses, such as scoring the responses or determining the presence of hallucinations in responses. Existing end-to-end evaluation methods can broadly be categorized into two types. The first type is LLM-based evaluation which implements an advanced LLM (e.g., GPT-4) to rate LVLM generated content based on hallucination Liu et al. (2023a); Sun et al. (2023). These methods leverage the LLM’s robust natural language understanding and processing capabilities. By integrating visual information (such as dense captions and object bounding boxes), user instructions, and model responses as input, LLMs are prompted to evaluate and score these responses. The second type is hallucination data driven model evaluation which constructs labelled hallucination datasets for fine-tuning models to detect hallucinations. Gunjal et al. (2023) creates the M-HalDetect dataset with annotated LVLM image descriptions and fine-tune an InstructBLIP model on the dataset for hallucination identification. Similarly, Wang et al. (2023c) compiles a dataset of hallucinatory responses generated by ChatGPT and then fine-tune a LLaMA model with LoRA Hu et al. (2022) on this dataset to differentiate hallucinatory responses from non-hallucinatory ones.
@@ 3.2 Evaluation on Hallucination Discrimination
Hallucination discrimination evaluation approach aim to assess the hallucination discrimination ability of LVLMs. The methods follow this approach typically adopt a question-answering format, inquiring LVLMs with descriptions that agree or conflict with the content of the provided input image and assessing the models’ responses Li et al. (2023b); Hu et al. (2023); Lovenia et al. (2023). POPE Li et al. (2023b) designs binary (Yes-or-No) questions about object presence in images such as “Is there a person in the image?” to evaluate the hallucination discrimination ability of LVLMs. The objects asked in questions are selected under three distinct sampling strategies: random (selecting random absent objects), popular (choosing frequently absent objects), and adversarial (selecting absent objects often co-occurring with present ones). CIEM Hu et al. (2023) is akin to POPE, yet it automates object selection by prompting ChatGPT. NOPE Lovenia et al. (2023) is another VQA-based method that is designed to evaluate LVLMs’ ability of discerning the absence of objects in visual queries, with correct responses being negative statements.
@@ 3.3 Evaluation Benchmarks
As delineated in Table 1, current research emphasizes generative benchmarks over discriminative ones. While discriminative benchmarks primarily evaluate hallucinations on object level, generative benchmarks extend their scope to include a broader spectrum of hallucinations, including attribute and relation ones Gunjal et al. (2023); Liu et al. (2023a); Jing et al. (2023); Wang et al. (2023c); Sun et al. (2023). Notably, AMBER Wang et al. (2023b) stands out as a comprehensive benchmark that integrates both generative and discriminative tasks. Furthermore, we observe that the metrics employed in generative benchmarks tend to be more complex and varied, in contrast to the reliance on accuracy metrics in discriminative benchmarks. This is probably due to that, distinct from discriminative methods that mostly compare model outputs with ground-truth answers, generative evaluation methods require tailored metrics to target specific hallucination categories in an attempt to analyze the responses generated by LVLMs.
@@ 4.2 Hallucinations from Vision Encoder
CLIP aligns visual content and word embeddings through contrastive learning that majorly focuses only on salient objects, falling to capture fine-grained aspects of an image and resulting in hallucinations Jain et al. (2023); Cho et al. (2022), such as background description, object counting, and object relation.
@@ 5.2 Mitigation for Vision Encoder
Most existing LVLMs employ the ViT from CLIP as the vision encoder, which majorly focuses only on salient objects and inevitably omit some visual cues. To enhance the object-level perceptual capability, Jain et al. (2023) proposes VCoder that uses extra perception modalities, such as segmentation map, depth map, or both, as the control inputs through additional vision encoders. To enhance the spatial awareness capability, Zhao et al. (2023a) suggests introducing addition pre-trained models to acquire spatial position information and scene graph details, which are then used to guide the LVLMs in addressing user queries.
@@ 5.3 Mitigation for Connection Module
Connection Modules Enhancing
@@ 7 Conclusion
Equipped with advanced vision encoders, robust LLMs, and modality alignment modules, LVLMs excel in open-domain vision-language tasks. However, hallucinations significantly challenge LVLMs’ practical application. In this survey, we conduct a meticulous investigation into the phenomenon of hallucinations in LVLMs. This exploration covers a detailed analysis of the fundamental causes behind these hallucinations, assesses innovative evaluation methods along with pertinent benchmarks, and discusses effective mitigation approaches. We also delve into the existing challenges and discuss potential directions. This survey aims to lay a foundation for tackle complexities of hallucinations in LVLMs and facilitate future research towards practical implementation of these models in various applications.

### v2

@@ Abstract
Recent development of Large Vision-Language Models (LVLMs) has attracted growing attention within the AI landscape for its practical implementation potential. However, “hallucination”, or more specifically, the misalignment between factual visual content and corresponding textual generation, poses a significant challenge of utilizing LVLMs. In this comprehensive survey, we dissect LVLM-related hallucinations in an attempt to establish an overview and facilitate future mitigation. Our scrutiny starts with a clarification of the concept of hallucinations in LVLMs, presenting a variety of hallucination symptoms and highlighting the unique challenges inherent in LVLM hallucinations. Subsequently, we outline the benchmarks and methodologies tailored specifically for evaluating hallucinations unique to LVLMs. Additionally, we delve into an investigation of the root causes of these hallucinations, encompassing insights from the training data and model components. We also critically review existing methods for mitigating hallucinations. The open questions and future directions pertaining to hallucinations within LVLMs are discussed to conclude this survey. To keep track of this field and continuously update our survey, we maintain a repository of relevant references at https://github.com/lhanchao777/LVLM-Hallucinations-Survey.
@@ 1 Introduction
Existing studies Rohrbach et al. (2018); Li et al. (2023b); Hu et al. (2023); Zhai et al. (2023) have addressed the issue of hallucination in image captioning models with a predominant focus on the “presence of objects”, specifically whether the objects depicted in a given image are accurately described by the text generated by the model. In contrast to image captioning models trained within closed domains, LVLMs leverage the robust understanding and expressive capabilities of LLMs to obtain more detailed and fluent generated descriptions. However, these enhanced abilities also diversify and potentially exacerbate hallucination, which is not only limited to object existence, but also manifested in the descriptive errors such as attribute and relation errors. We focus on visual hallucinations, referring to all inconsistencies between the semantic content conveyed by the image and the textual content generated by the model.
@@
While the causes of hallucinations in LLMs Huang et al. (2023a) have been discussed extensively, LVLMs’ visual modality introduces unique challenges in analyzing these occurrences. We conduct a thorough analysis of hallucinations in LVLMs, focusing on the training data and model characteristics. Our analysis indicates that hallucinations in LVLMs are caused not only by the generative nature of LLMs but also by biased training data, the inability of vision encoders to accurately ground images, misalignment among different modalities, insufficient context attention, and many other factors. Following this, we provide a comprehensive overview of the existing hallucination mitigation methods. Corresponding to the causes, current mitigation approaches predominantly focus on the optimization of training data, refinement of various modules within LVLMs, and post-processing of generated outputs. These methods are employed to diminish the occurrence of hallucinations, thereby yielding more faithful responses. In the end, we list several considerable directions for developing the research of hallucinations in LVLMs.
@@ 2.1 Large Vision-Language Models
LVLM architectures typically comprise three components: a visual encoder, a modality connection module, and a LLM. The visual encoder, often an adaptation of the CLIP vision encoder Radford et al. (2021), transforms input images into visual tokens. The connection module is designed to align visual tokens with the word embedding space of the LLM, ensuring that the LLM can handle the visual information. There are various methods for modality alignment, including cross-attention Alayrac et al. (2022), adapters Gao et al. (2023), Q-Formers Li et al. (2023a); Dai et al. (2023a); Zhu et al. (2023), and simpler structures such as linear layers or multi-layer perceptrons (MLP) Liu et al. (2023c); Chen et al. (2023b); Liu et al. (2023b). The LLM functions like the central processing unit within LVLMs, receiving aligned visual and textual information and subsequently synthesizing this information to produce responses.
@@ 3 Evaluation Methods and Benchmarks
FAITHSCORE Jing et al. (2023) , leaf, text width=15em ] ] [ Model-based
@@ 3.1 Evaluation on Non-Hallucinatory Generation
End-to-end methods directly assess the performance of LVLMs by evaluating their responses, such as scoring the responses or determining the presence of hallucinations in responses. Existing end-to-end evaluation methods can broadly be categorized into two types. The first type is LLM-based evaluation which implements an advanced LLM (e.g., GPT-4) to rate LVLM generated content based on hallucination Liu et al. (2023a); Sun et al. (2023). These methods leverage the LLM’s robust natural language understanding and processing capabilities. By integrating visual information (such as dense captions and object bounding boxes), user instructions, and model responses as input, LLMs are prompted to evaluate and score these responses. The second type is hallucination data driven model evaluation which constructs labelled hallucination datasets for fine-tuning models to detect hallucinations. Gunjal et al. (2023) creates the M-HalDetect dataset with annotated LVLM image descriptions and fine-tune an InstructBLIP Dai et al. (2023a) model on the dataset for hallucination identification. Similarly, Wang et al. (2023c) compiles a dataset of hallucinatory responses generated by ChatGPT and then fine-tune a LLaMA model with LoRA Hu et al. (2022) on this dataset to differentiate hallucinatory responses from non-hallucinatory ones.
@@ 3.2 Evaluation on Hallucination Discrimination
Hallucination discrimination evaluation approach aim to assess the hallucination discrimination ability of LVLMs. The methods follow this approach typically adopt a question-answering format, posing inquiries to LVLMs consisting of descriptions that agree or conflict with the content of the provided input image and assessing the models’ responses Li et al. (2023b); Hu et al. (2023); Lovenia et al. (2023). POPE Li et al. (2023b) designs binary (Yes-or-No) questions about object presence in images such as “Is there a person in the image?” to evaluate the hallucination discrimination ability of LVLMs. The objects asked in questions are selected under three distinct sampling strategies: random (selecting random absent objects), popular (choosing most frequent objects in dataset but absent in current image), and adversarial (selecting absent objects often co-occurring with present ones). CIEM Hu et al. (2023) is akin to POPE, yet it automates object selection by prompting ChatGPT. NOPE Lovenia et al. (2023) is another VQA-based method that is designed to evaluate LVLMs’ ability of discerning the absence of objects in visual queries, with correct responses being negative statements.
@@ 3.3 Evaluation Benchmarks
As delineated in Table 1, current research emphasizes generative benchmarks over discriminative ones. While discriminative benchmarks primarily evaluate hallucinations at the object level, generative benchmarks extend their scope to include a broader spectrum of hallucinations, including attribute and relation ones Gunjal et al. (2023); Liu et al. (2023a); Jing et al. (2023); Wang et al. (2023c); Sun et al. (2023). Notably, AMBER Wang et al. (2023b) stands out as a comprehensive benchmark that integrates both generative and discriminative tasks. Furthermore, we observe that the metrics employed in generative benchmarks tend to be more complex and varied, in contrast to the reliance on accuracy metrics in discriminative benchmarks. This is probably because, distinct from discriminative methods that mostly compare model outputs with ground-truth answers, generative evaluation methods require tailored metrics to target specific hallucination categories in an attempt to analyze the responses generated by LVLMs.
@@ 4.2 Hallucinations from Vision Encoder
CLIP aligns visual content and word embeddings through contrastive learning that mainly focuses only on salient objects, failing to capture fine-grained aspects of an image and resulting in hallucinations Jain et al. (2023); Cho et al. (2022), such as background description, object counting, and object relation.
@@ 5.2 Mitigation for Vision Encoder
Most existing LVLMs employ the ViT from CLIP as the vision encoder, which mainly focuses only on salient objects and inevitably omit some visual cues. To enhance the object-level perceptual capability, Jain et al. (2023) proposes VCoder that uses extra perception modalities, such as segmentation map, depth map, or both, as the control inputs through additional vision encoders. To enhance the spatial awareness capability, Zhao et al. (2023a) suggests introducing addition pre-trained models to acquire spatial position information and scene graph details, which are then used to guide the LVLMs in addressing user queries.
@@ 5.3 Mitigation for Connection Module
Connection Module Enhancing
@@ 7 Conclusion
Equipped with advanced vision encoders, robust LLMs, and modality connection modules, LVLMs excel in open-domain vision-language tasks. However, hallucinations significantly challenge LVLMs’ practical application. In this survey, we conduct a meticulous investigation into the phenomenon of hallucinations in LVLMs. This exploration introduces innovative evaluation methods along with pertinent benchmarks, provides a detailed analysis of the fundamental causes behind these hallucinations, and discusses effective mitigation approaches. We also delve into the existing challenges and discuss potential directions. This survey aims to lay a foundation for tackle complexities of hallucinations in LVLMs and facilitate future research towards practical implementation of these models in various applications.